# Hybrid Generative Fusion for Efficient and Privacy-Preserving Face Recognition Dataset Generation

### v1

@@ 1 Introduction
Recent advances in face recognition (FR) have been largely driven by deep learning models [23, 11] trained on large-scale, labeled datasets. However, due to growing privacy, ethical, and legal concerns, the use of real facial identities  [22, 26, 18] is increasingly restricted or infeasible. The DataCV ICCV Face Recognition Dataset Construction Challenge poses this issue by introducing the task of creating high-quality training datasets [27, 13, 6, 7, 21] that do not include any real-world identities. The objective is to train a fixed face recognition model with only the synthetic identities, aiming to achieve performance on test sets that is comparable to or better than models trained on real identities.
@@
    Intra-Class Variation: For robust generalization, FR models must handle variations [19, 5, 24, 25, 10, 16] in pose, lighting, expression, and background. Synthetic datasets often lack this natural diversity, especially when samples are generated from a single reference image or through narrowly constrained pipelines.
@@
To make up for the vacancy after the cleaning and ensure diversity, we generate new identities using Stable Diffusion [20] guided by prompt engineering [15, 12]. For each generated identity, a single reference image is created, and 49 additional samples are synthesized using Vec2Face [27], resulting in visually consistent image sets. However, these generated samples tend to be overly similar. To mitigate the risk of overfitting, we apply a curriculum learning [4] strategy, positioning these easy (low intra-class variance) identities at the beginning of the training dataset, followed by harder (high intra-class variance) identities [2, 3] from the cleaned HSFace set. This gradual exposure improves the model’s ability to generalize.

### v2

@@ 1 Introduction
Recent advances in face recognition (FR) have been largely driven by deep learning models [23, 11] trained on large-scale, labeled datasets. However, due to growing privacy, ethical, and legal concerns, the use of real facial identities  [22, 26, 18, 33, 30] is increasingly restricted or infeasible. The DataCV ICCV Face Recognition Dataset Construction Challenge poses this issue by introducing the task of creating high-quality training datasets [31, 13, 6, 7, 21, 34, 29, 28] that do not include any real-world identities. The objective is to train a fixed face recognition model with only the synthetic identities, aiming to achieve performance on test sets that is comparable to or better than models trained on real identities.
@@
    Intra-Class Variation: For robust generalization, FR models must handle variations [19, 5, 24, 25, 10, 16, 27] in pose, lighting, expression, and background. Synthetic datasets often lack this natural diversity, especially when samples are generated from a single reference image or through narrowly constrained pipelines.
@@
To make up for the vacancy after the cleaning and ensure diversity, we generate new identities using Stable Diffusion [20] guided by prompt engineering [15, 12, 35]. For each generated identity, a single reference image is created, and 49 additional samples are synthesized using Vec2Face [31], resulting in visually consistent image sets. However, these generated samples tend to be overly similar. To mitigate the risk of overfitting, we apply a curriculum learning [4, 32] strategy, positioning these easy (low intra-class variance) identities at the beginning of the training dataset, followed by harder (high intra-class variance) identities [2, 3, 33, 36] from the cleaned HSFace set. This gradual exposure improves the model’s ability to generalize.